OpenStudy (anonymous):

Calculus question! lim f(θ) where f(θ) = {3cos(2θ ) if θ ≤ π θ→ π+ { (sin(θ/2)) / (θ) if θ>π Can anyone show me a step by step solution how to do this so I know how to do the other ones similar to it? Please and thank you! All I know is that I am supposed to use the (sin(θ/2)) / (θ) if θ>π equation. So what do I do next?

OpenStudy (anonymous):

You want to do the left hand and right hand limits.

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

\[ \lim_{\theta\to \pi^+}3\cos(2\theta) = 3\cos(2\pi) = 3 \]

OpenStudy (anonymous):

\[ \lim_{\theta\to \pi^-}\frac{\sin(\theta/2)}{\theta} = \frac{\sin(\pi/2)}{\pi} = \frac 1 \pi \]

OpenStudy (anonymous):

\[ \lim_{\theta\to\pi^-}f(\theta) \neq \lim_{\theta\to\pi^+}f(\theta) \]The limit does not exist.

OpenStudy (anonymous):

Oh! I see, now it makes sense! Thank you so much! SO because the left and right hand sides are not equal to each other, the limit does not exist?

OpenStudy (anonymous):

That is a basic fact about limits.